# Blessing of dimension in Bayesian inference on covariance matrices

### v1

@@ Acknowledgments
This work was partially supported by Merck, the United States Office of Naval Research (ONR) under grant N00014-21-1-2510, the European Research Council (ERC) under grant 856506, and the National Institutes of Health (NIH) under grant R01ES035625. Code used to implement CC-FABLE used in the simulation experiments and the gene expression data application is available to download at https://github.com/shounakch/FABLE.
@@
The gene expression data set used in Section 5 of this paper is publicly available at https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE109125.
@@
    Section F contains the proof of relevant lemmas used to prove Theorems 3.7 and 3.8 in the main paper.
@@
    For any $\widehat{\mathbf{M}}$ satisfying the condition, $\mathbf{K}=\{\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+(\mathbb{I}_{k}/\tau^{2})\}^{-1}=(n+\tau^{-2})^{-1}\mathbb{I}_{k}.$ It is clear that the pseudo-posterior of $\widetilde{\sigma}_{j}^{2}$ only depends on $\widehat{\mathbf{M}}$ through $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}$ for all $1\leq j\leq p$. For $1\leq u\leq p,$ a pseudo-posterior draw of $\tilde{\lambda}_{u}$ may be represented as
    $\widetilde{\lambda}_{u}=\mu_{u}+\widetilde{\sigma}_{u}\mathbf{K}^{1/2}Z_{u},$
@@
    $\widetilde{\lambda}_{u}^{\top}\widetilde{\lambda}_{v}=\mu_{u}^{\top}\mu_{v}+\widetilde{\sigma}_{v}\mu_{u}^{\top}\mathbf{K}^{1/2}Z_{v}+\widetilde{\sigma}_{u}\mu_{v}^{\top}\mathbf{K}^{1/2}Z_{u}+\widetilde{\sigma}_{u}\widetilde{\sigma}_{v}Z_{u}^{\top}\mathbf{K}Z_{v}.$
    We now consider each expression one-by-one. The first term is $\mu_{u}^{\top}\mu_{v}=y^{(u)\top}\widehat{\mathbf{M}}\mathbf{K}^{2}\widehat{\mathbf{M}}^{\top}y^{(v)}=(n+\tau^{-2})^{-2}y^{(u)\top}\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}y^{(v)}.$ When $u\neq v$, the second and third terms have Gaussian distributions with means 00 and variances $(n+\tau^{-2})^{-1}\widetilde{\sigma}_{u}^{2}\|\mu_{v}\|_{2}^{2}$ and $(n+\tau^{-2})^{-1}\widetilde{\sigma}_{v}^{2}\|\mu_{u}\|_{2}^{2}$, respectively. When $u=v$, both terms are equal and the sum has a Gaussian distribution with mean 00 and variance $4\widetilde{\sigma}_{u}^{2}\|\mu_{u}\|_{2}^{2}$. The final term is equal to $(n+\tau^{-2})^{-1}\widetilde{\sigma}_{u}\widetilde{\sigma}_{v}Z_{u}^{\top}Z_{v}$. Thus, all the terms depend on $\widehat{\mathbf{M}}$ only as hypothesized, implying the result for $\widetilde{L}$. The result for $\widetilde{\Psi}$ is obtained by simply observing that $\widetilde{\Psi}=\widetilde{L}+\widetilde{\Sigma}.$
@@ B.1 Proof of part (a)
We first start with the pseudo-posterior contraction of $L_{C}$ to $L_{0}.$ Let $G_{0}=\sqrt{n}\mathbf{Y}^{\top}U/(n+\tau^{-2})$ and $G=G_{0}G_{0}^{\top}=n\mathbf{Y}^{\top}UU^{\top}\mathbf{Y}/(n+\tau^{-2})^{2}.$ We have
$\begin{split}L_{C}-L_{0}&={\Lambda}_{C}{\Lambda}_{C}^{\top}-\Lambda_{0}\Lambda_{0}^{\top}\\ &=\left[\frac{\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}+\widetilde{E}\widehat{M}^{\top}\mathbf{Y}}{n+\tau^{-2}}+\widetilde{E}\widetilde{E}^{\top}\right]+\left(G-\Lambda_{0}\Lambda_{0}^{\top}\right),\end{split}$
by expressing $\Lambda_{C}=G_{0}+\widetilde{E}$, where $\widetilde{E}=[\tilde{e}_{1},\ldots,\tilde{e}_{p}]^{\top}$ with $\tilde{e}_{j}\overset{ind}{\sim}N_{k}\left(0,\frac{\rho^{2}\sigma_{C,j}^{2}}{n+\tau^{-2}}\mathbb{I}_{k}\right)$ for $j=1,\ldots,p$. We decompose
@@
To develop an upper bound of $\|L_{C}-L_{0}\|$, we aim to develop an upper bound for the spectral norm of each term in the inequality above. We enumerate them as follows.
@@
    First, we develop a probabilistic upper bound for ${\sigma}_{C,j}^{2}.$ Recall that $\|y^{(j)}\|_{2}^{2}=\sum_{i=1}^{n}y_{ij}^{2}\sim(\sigma_{0j}^{2}+\|\lambda_{0j}\|_{2}^{2})\,\chi_{n}^{2}.$ We apply the tail inequality bound of the $\chi^{2}$ distribution following Lemma 1 in Laurent and Massart, (2000) to obtain $\|y^{(j)}\|_{2}\lesssim n$ with probability at least $1-o(1)$. Thus, with probability at least $1-o(1)$ we have
@@
    ${\sigma}_{C,j}^{2}\sim\mbox{IG}\left(\frac{\gamma_{0}+n}{2},\frac{\gamma_{0}\delta_{0}^{2}+\|y^{(j)}\|_{2}^{2}-(n+\tau^{-2})\|\mu_{j}\|_{2}^{2}}{2}\right),$
@@
    $\begin{split}\Pi_{C}\left(\max_{1\leq j\leq p}{\sigma}_{C,j}^{2}\leq C^{\prime}\right)=1-o_{P_{0}}(1).\end{split}$ (20)
@@
    Recall $M_{0}$ is an $n$-by-$k$ matrix with i.i.d. Gaussian entries. By Lemma E.1, we have with probability at least $1-o(1)$ the following:
@@
    $\lesssim\frac{p_{n}(k+\sqrt{nk})}{n}\lesssim\dfrac{p_{n}}{\sqrt{n}}.$
@@
    By Lemma E.2, $\|E^{\top}E\|=\|E\|^{2}\lesssim n+p_{n}+\sqrt{np_{n}}$ with probability at least $1-o(1)$ and $\|\Lambda_{0}M_{0}^{\top}E\|\leq\|\Lambda_{0}\|\,\|M_{0}\|\,\|E\|\lesssim\sqrt{np_{n}}(\sqrt{n}+\sqrt{p_{n}})$ with probability at least $1-o(1)$. Thus, with probability at least $1-o(1)$,
@@
    $\begin{split}\left\|\frac{\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}+\widetilde{E}\widehat{M}^{\top}\mathbf{Y}}{n+\tau^{-2}}+\widetilde{E}\widetilde{E}^{\top}\right\|\lesssim\dfrac{2}{n}\left\|\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}\right\|+\|\widetilde{E}\widetilde{E}^{\top}\|.\end{split}$
@@
    $\|\widetilde{E}\widetilde{E}^{\top}\|=\|\widetilde{E}\|^{2}\lesssim\frac{p_{n}}{n}\rho^{2}\left(\max_{1\leq j\leq p}{\sigma}_{C,j}^{2}\right)\lesssim\frac{p_{n}}{n}.$
    $\frac{1}{n}\left\|\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}\right\|$ $\lesssim\frac{1}{n}\|\mathbf{Y}^{\top}\sqrt{n}U\|\left\|\widetilde{E}\right\|\lesssim\frac{\rho}{\sqrt{n}}\|\mathbf{Y}\|\sqrt{\frac{p_{n}}{n}\left(\max_{1\leq j\leq p}{\sigma}^{2}_{C,j}\right)}$
@@
    $\left\|\frac{\mathbf{Y}^{\top}\widehat{M}\widetilde{E}^{\top}+\widetilde{E}\widehat{M}^{\top}\mathbf{Y}}{(n+\tau^{-2})}+\widetilde{E}\widetilde{E}^{\top}\right\|\lesssim\dfrac{p_{n}}{\sqrt{n}}.$
@@ B.2 Proof of part (b)
We now show the contraction result for the pseudo-posterior of $\Sigma_{C}$. We start out by observing that $||{\Sigma}_{C}-\Sigma_{0}||=\max_{1\leq j\leq p}|D_{j}|$, where $D_{j}={\sigma}_{C,j}^{2}-\sigma_{0j}^{2}$ for $j=1,\ldots,p.$ Let $\kappa_{j}={\sigma}_{C,j}^{-2}$, so that $\kappa_{j}\overset{ind}{\sim}G(\gamma_{n}/2,\gamma_{n}\delta_{j}^{2}/2)$ under $\Pi_{C}$. Let
$U_{j}=\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\kappa_{j}-\dfrac{\gamma_{n}}{2}.$
@@
$D_{j}=\left(1+\dfrac{2U_{j}}{\gamma_{n}}\right)^{-1}\,\left\{(\delta_{j}^{2}-\sigma_{0j}^{2})-\dfrac{2U_{j}}{\gamma_{n}}\sigma_{0j}^{2}\right\}.$
Using Lemma E.7, we have $\max_{1\leq j\leq p}|U_{j}|/\gamma_{n}\lesssim\{(\log p_{n})/n\}^{1/3}$ with probability at least $1-o_{P_{0}}(1).$ Thus, $\min_{1\leq j\leq p}|1+(2U_{j}/\gamma_{n})|\gtrsim 1/2$ with probability at least $1-o_{P_{0}}(1).$ Therefore with probability at least $1-o_{P_{0}}(1),$
$\max_{1\leq j\leq p}|D_{j}|\lesssim 2\max_{1\leq j\leq p}|\delta_{j}^{2}-\sigma_{0j}^{2}|+4\sigma_{0j}^{2}\max_{1\leq j\leq p}\dfrac{|U_{j}|}{\gamma_{n}}.$
@@
where $Z_{j}/\sigma_{0j}^{2}\sim\chi_{n-k}^{2}\equiv G\{(n-k)/2,1/2\}$ and $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}$ with probability at least $1-o(1).$ Using Lemma E.7, we obtain
@@
with probability at least $1-o(1)$. Thus, with probability at least $1-o(1)$, we get
@@
From Lemma E.7, we have $\max_{1\leq j\leq p}|U_{j}|/\gamma_{n}\lesssim\left({\log p_{n}}/{n}\right)^{1/3}$ with probability at least $1-o_{P_{0}}(1)$. Thus, we obtain
@@
$\|{\Sigma}_{C}-\Sigma_{0}\|\lesssim\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}$
@@ B.3 Proof of part (c)
Under the conditions of Theorem 1, we have $\|\Psi_{0}\|\asymp\|L_{0}\|\asymp p_{n}.$ By the triangle inequality, $\|\Psi_{C}-\Psi_{0}\|\leq\|L_{C}-L_{0}\|+\|\Sigma_{C}-\Sigma_{0}\|$. Thus, we have under $\Pi_{C}$ that
$\dfrac{\|{\Psi}_{C}-\Psi_{0}\|}{\|\Psi_{0}\|}\lesssim\dfrac{1}{\sqrt{p_{n}}}+\dfrac{1}{\sqrt{n}}+o\left(\dfrac{1}{p_{n}}\right)\lesssim\dfrac{1}{\sqrt{p_{n}}}+\dfrac{1}{\sqrt{n}}$
with probability at least $1-o_{P_{0}}(1)$, proving the result. ∎
@@
First let $u\neq v$, so that $\Psi_{C,uv}=L_{C,uv}$, where $L_{C,uv}$ is the $uv$th element of $L_{C}=\Lambda_{C}\Lambda_{C}^{\top}$. According to the pseudo-posterior generation mechanism after coverage correction,
$L_{C,uv}=\mu_{u}^{\top}\mu_{v}+\dfrac{\rho\left({\sigma}_{C,v}\mu_{u}^{\top}\tilde{e}_{C,v}+{\sigma}_{C,u}\mu_{v}^{\top}\tilde{e}_{C,u}\right)}{\sqrt{n+\tau^{-2}}}+\dfrac{\rho^{2}{\sigma}_{C,u}{\sigma}_{C,v}\tilde{e}_{C,u}^{\top}\tilde{e}_{C,v}}{n+\tau^{-2}},$
for $\widetilde{e}_{C,u},\widetilde{e}_{C,v}\overset{ind}{\sim}N_{k}(0,\mathbb{I}_{k})$. We can express
$\rho\left(\sigma_{C,v}\mu_{u}^{\top}\tilde{e}_{C,v}+\sigma_{C,u}\mu_{v}^{\top}\tilde{e}_{C,u}\right)$ $=l_{0,uv}(\rho)R_{uv}+d_{uv}(\rho)R_{uv},$
for $R_{uv}\sim N(0,1)$ such that $R_{uv}$ is independent of $\mathbf{Y},\sigma_{C,u}^{2},$ and $\sigma_{C,v}^{2}$, with $l_{0,uv}^{2}(\rho)=\rho^{2}\left(\sigma_{0v}^{2}\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2}\|\lambda_{0v}\|_{2}^{2}\right)$ and $d_{uv}(\rho)=\rho(\sigma_{C,v}^{2}||\mu_{u}||_{2}^{2}+\sigma_{C,u}^{2}||\mu_{v}||_{2}^{2})^{1/2}-l_{0,uv}(\rho)$. We can now express
$\sqrt{n}(L_{C,uv}-\mu_{u}^{\top}\mu_{v})=\dfrac{\sqrt{n}}{\sqrt{n+\tau^{-2}}}l_{0,uv}(\rho)R_{uv}+d_{uv}^{*}(\rho),$

### v3

@@ Acknowledgments
This work was partially supported by Merck, the United States Office of Naval Research (ONR) under grant N00014-21-1-2510, the European Research Council (ERC) under grant 856506, and the National Institutes of Health (NIH) under grant R01ES035625. Code used to implement FABLE in the simulation experiments and the gene expression data application is available to download at https://github.com/shounakch/FABLE.
@@
The gene expression dataset used in Section 5 of this paper is publicly available for download at https://www.ncbi.nlm.nih.gov/geo/query/acc.cgi?acc=GSE109125 with filename GSE109125_Genes_count_table.tsv.gz.
@@
    Section F contains the proof of relevant lemmas used to prove Theorems 3.8 and 3.9 in the main paper.
  7. 7.
    Section G contains additional simulation results for cases described in the main paper.
  8. 8.
    Section H contains information regarding the gene data application described in the main paper.
  9. 9.
    Section I contains details regarding the FABLE algorithm.
@@
    For any $\widehat{\mathbf{M}}$ satisfying the condition, $\mathbf{K}=\{\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}+(\mathbb{I}_{k}/\tau^{2})\}^{-1}=(n+\tau^{-2})^{-1}\mathbb{I}_{k}.$ It is clear that the FABLE-posterior of $\widetilde{\sigma}_{j}^{2}$ only depends on $\widehat{\mathbf{M}}$ through $\widehat{\mathbf{M}}^{\top}\widehat{\mathbf{M}}$ and $\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}$ for all $1\leq j\leq p$. For $1\leq u\leq p,$ a FABLE-posterior draw of $\widetilde{\lambda}_{u}$ may be represented as
    $\widetilde{\lambda}_{u}=\mu_{u}+\rho\,\widetilde{\sigma}_{u}\mathbf{K}^{1/2}Z_{u},$
@@
    $\widetilde{\lambda}_{u}^{\top}\widetilde{\lambda}_{v}=\mu_{u}^{\top}\mu_{v}+\rho\,\widetilde{\sigma}_{v}\mu_{u}^{\top}\mathbf{K}^{1/2}Z_{v}+\rho\,\widetilde{\sigma}_{u}\mu_{v}^{\top}\mathbf{K}^{1/2}Z_{u}+\rho^{2}\,\widetilde{\sigma}_{u}\widetilde{\sigma}_{v}Z_{u}^{\top}\mathbf{K}Z_{v}.$
    We now consider each expression one-by-one. The first term is $\mu_{u}^{\top}\mu_{v}=y^{(u)\top}\widehat{\mathbf{M}}\mathbf{K}^{2}\widehat{\mathbf{M}}^{\top}y^{(v)}=(n+\tau^{-2})^{-2}y^{(u)\top}\widehat{\mathbf{M}}\widehat{\mathbf{M}}^{\top}y^{(v)}.$ When $u\neq v$, the second and third terms have Gaussian distributions with means 0 and variances $\rho^{2}\,\widetilde{\sigma}_{u}^{2}\|\mu_{v}\|_{2}^{2}/(n+\tau^{-2})$ and $\rho^{2}\,\widetilde{\sigma}_{v}^{2}\|\mu_{u}\|_{2}^{2}/(n+\tau^{-2})$, respectively. When $u=v$, both terms are equal and the sum has a Gaussian distribution with mean 0 and variance $4\widetilde{\sigma}_{u}^{2}\|\mu_{u}\|_{2}^{2}$. The final term is equal to $(n+\tau^{-2})^{-1}\widetilde{\sigma}_{u}\widetilde{\sigma}_{v}Z_{u}^{\top}Z_{v}$. Thus, all the terms depend on $\widehat{\mathbf{M}}$ only as hypothesized, implying the result for $\widetilde{L}$. The result for $\widetilde{\Psi}$ is obtained by simply observing that $\widetilde{\Psi}=\widetilde{L}+\widetilde{\Sigma}$.
@@ B.1 Proof of part (a)
We first start with the FABLE-posterior contraction of $\widetilde{L}$ to $L_{0}.$ Let $G_{0}=\sqrt{n}\mathbf{Y}^{\top}U/(n+\tau^{-2})$ and $G=G_{0}G_{0}^{\top}=n\mathbf{Y}^{\top}UU^{\top}\mathbf{Y}/(n+\tau^{-2})^{2}.$ We have
$\begin{split}\widetilde{L}-L_{0}&=\widetilde{\Lambda}\widetilde{\Lambda}^{\top}-\Lambda_{0}\Lambda_{0}^{\top}\\ &=\left[\frac{\mathbf{Y}^{\top}\widehat{\mathbf{M}}\widetilde{E}^{\top}+\widetilde{E}\widehat{\mathbf{M}}^{\top}\mathbf{Y}}{n+\tau^{-2}}+\widetilde{E}\widetilde{E}^{\top}\right]+\left(G-\Lambda_{0}\Lambda_{0}^{\top}\right),\end{split}$
by expressing $\widetilde{\Lambda}=G_{0}+\widetilde{E}$, where $\widetilde{E}=[\tilde{e}_{1},\ldots,\tilde{e}_{p}]^{\top}$ with $\tilde{e}_{j}\overset{ind}{\sim}N_{k}\left(0,\frac{\rho^{2}\widetilde{\sigma}_{j}^{2}}{n+\tau^{-2}}\mathbb{I}_{k}\right)$ for $j=1,\ldots,p$. We decompose
@@
To develop an upper bound of $\|\widetilde{L}-L_{0}\|$, we aim to develop an upper bound for the spectral norm of each term in the expression above. We enumerate them as follows.
@@
    First, we develop a probabilistic upper bound for $\widetilde{\sigma}_{j}^{2}.$ Recall that $\|y^{(j)}\|_{2}^{2}=\sum_{i=1}^{n}y_{ij}^{2}\sim(\sigma_{0j}^{2}+\|\lambda_{0j}\|_{2}^{2})\,\chi_{n}^{2}.$ We apply the tail inequality bound of the $\chi^{2}$ distribution following Lemma 1 in Laurent and Massart, (2000) to obtain $\|y^{(j)}\|_{2}^{2}\lesssim n$ with probability at least 1o(1)1-o(1). Thus, with probability at least 1o(1)1-o(1) we have
@@
    $\widetilde{\sigma}_{j}^{2}\sim\mbox{IG}\left(\frac{\gamma_{0}+n}{2},\frac{\gamma_{0}\delta_{0}^{2}+\|y^{(j)}\|_{2}^{2}-(n+\tau^{-2})\|\mu_{j}\|_{2}^{2}}{2}\right),$
@@
    $\begin{split}\Pi_{C}\left(\max_{1\leq j\leq p}\widetilde{\sigma}_{j}^{2}\leq C^{\prime}\right)=1-o_{P_{0}}(1).\end{split}$ (22)
@@
    Recall $M_{0}$ is an $n$-by-$k$ matrix with i.i.d. Gaussian entries. By Lemma E.1, we have with probability at least 1o(1)1-o(1) the following:
@@
    $\lesssim\frac{p_{n}\,\sqrt{n\log n}}{n}\lesssim p_{n}\sqrt{\dfrac{\log n}{n}}.$
@@
    By Lemma E.2, $\|E^{\top}E\|=\|E\|^{2}\lesssim n+p_{n}+\sqrt{np_{n}}$ with probability at least 1o(1)1-o(1) and $\|\Lambda_{0}M_{0}^{\top}E\|\leq\|\Lambda_{0}\|\,\|M_{0}\|\,\|E\|\lesssim\sqrt{np_{n}}(\sqrt{n}+\sqrt{p_{n}})$ with probability at least 1o(1)1-o(1). Thus, with probability at least 1o(1)1-o(1),
@@
[…]
@@
    $\|\widetilde{E}\widetilde{E}^{\top}\|=\|\widetilde{E}\|^{2}\lesssim\frac{p_{n}}{n}\rho^{2}\left(\max_{1\leq j\leq p}\widetilde{\sigma}_{j}^{2}\right)\lesssim\frac{p_{n}}{n}.$
    $\frac{1}{n}\left\|\mathbf{Y}^{\top}\widehat{\mathbf{M}}\widetilde{E}^{\top}\right\|$ $\lesssim\frac{1}{n}\|\mathbf{Y}^{\top}\sqrt{n}U\|\left\|\widetilde{E}\right\|\lesssim\frac{\rho}{\sqrt{n}}\|\mathbf{Y}\|\sqrt{\frac{p_{n}}{n}\left(\max_{1\leq j\leq p}\widetilde{\sigma}^{2}_{j}\right)}$
@@
    $\left\|\frac{\mathbf{Y}^{\top}\widehat{\mathbf{M}}\widetilde{E}^{\top}+\widetilde{E}\widehat{\mathbf{M}}^{\top}\mathbf{Y}}{(n+\tau^{-2})}+\widetilde{E}\widetilde{E}^{\top}\right\|\lesssim\dfrac{p_{n}}{\sqrt{n}}.$
@@ B.2 Proof of part (b)
We now show the contraction result for the FABLE-posterior of $\widetilde{\Sigma}$. We start out by observing that $\widetilde{\Sigma}-\Sigma_{0}||=\max_{1\leq j\leq p}|D_{j}|$, where $D_{j}=\widetilde{\sigma}_{j}^{2}-\sigma_{0j}^{2}$ for $j=1,\ldots,p.$ Let $\kappa_{j}=\widetilde{\sigma}_{j}^{-2}$, so that $\kappa_{j}\overset{ind}{\sim}G(\gamma_{n}/2,\gamma_{n}\delta_{j}^{2}/2)$ under $\widetilde{\Pi}$. Let
$\Delta_{j}=\dfrac{\gamma_{n}\delta_{j}^{2}}{2}\kappa_{j}-\dfrac{\gamma_{n}}{2}.$
@@
$D_{j}=\left(1+\dfrac{2\Delta_{j}}{\gamma_{n}}\right)^{-1}\,\left\{(\delta_{j}^{2}-\sigma_{0j}^{2})-\dfrac{2\Delta_{j}}{\gamma_{n}}\sigma_{0j}^{2}\right\}.$
Using Lemma E.7, we have $\max_{1\leq j\leq p}|\Delta_{j}|/\gamma_{n}\lesssim\{(\log p_{n})/n\}^{1/3}$ with probability at least $1-o_{P_{0}}(1).$ Thus, $\min_{1\leq j\leq p}|1+(2\Delta_{j}/\gamma_{n})|\gtrsim 1/2$ with probability at least $1-o_{P_{0}}(1).$ Therefore with probability at least $1-o_{P_{0}}(1),$
$\max_{1\leq j\leq p}|D_{j}|\lesssim 2\max_{1\leq j\leq p}|\delta_{j}^{2}-\sigma_{0j}^{2}|+4\sigma_{0j}^{2}\max_{1\leq j\leq p}\dfrac{|\Delta_{j}|}{\gamma_{n}}.$
@@
where $Z_{j}/\sigma_{0j}^{2}\sim\chi_{n-k}^{2}\equiv G\{(n-k)/2,1/2\}$ and $\max_{1\leq j\leq p}|F_{j}|\lesssim\dfrac{1}{n}+\dfrac{1}{p_{n}}$ with probability at least 1o(1).1-o(1). Using Lemma E.7, we obtain
@@
with probability at least 1o(1)1-o(1). Thus, with probability at least 1o(1)1-o(1), we get
@@
From Lemma E.7, we have $\max_{1\leq j\leq p}|\Delta_{j}|/\gamma_{n}\lesssim\left({\log p_{n}}/{n}\right)^{1/3}$ with probability at least $1-o_{P_{0}}(1)$. Thus, we obtain
@@
$\|\widetilde{\Sigma}-\Sigma_{0}\|\lesssim\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}$
@@ B.3 Proof of part (c)
Under the conditions of Theorem 1, we have $\|\Psi_{0}\|\asymp\|L_{0}\|\asymp p_{n}.$ By the triangle inequality, $\|\widetilde{\Psi}-\Psi_{0}\|\leq\|\widetilde{L}-L_{0}\|+\|\widetilde{\Sigma}-\Sigma_{0}\|$. Thus, under $\widetilde{\Pi}$, we have
$\dfrac{\|\widetilde{\Psi}-\Psi_{0}\|}{\|\Psi_{0}\|}\lesssim\dfrac{1}{\sqrt{p_{n}}}+\sqrt{\dfrac{\log n}{n}}+\dfrac{1}{p_{n}}\left[\left(\dfrac{\log p_{n}}{n}\right)^{1/3}+\dfrac{1}{p_{n}}\right]\lesssim\dfrac{1}{\sqrt{p_{n}}}+\sqrt{\dfrac{\log n}{n}}$
with probability at least $1-o_{P_{0}}(1)$, using the results of parts (a) and (b). This proves the result. ∎
@@
First let $u\neq v$, so that $\widetilde{\Psi}_{uv}=\widetilde{L}_{uv}$, where $\widetilde{L}_{uv}$ is the $uv$th element of $\widetilde{L}=\widetilde{\Lambda}\widetilde{\Lambda}^{\top}$. According to the FABLE-posterior generation mechanism after coverage correction,
$\widetilde{L}_{uv}=\mu_{u}^{\top}\mu_{v}+\dfrac{\rho\left(\widetilde{\sigma}_{v}\mu_{u}^{\top}\tilde{e}_{v}+\widetilde{\sigma}_{u}\mu_{v}^{\top}\tilde{e}_{u}\right)}{\sqrt{n+\tau^{-2}}}+\dfrac{\rho^{2}\widetilde{\sigma}_{u}\widetilde{\sigma}_{v}\tilde{e}_{u}^{\top}\tilde{e}_{v}}{n+\tau^{-2}},$
for $\tilde{e}_{u},\tilde{e}_{v}\overset{ind}{\sim}N_{k}(0,\mathbb{I}_{k})$. Thus, we can express
$\rho\left(\widetilde{\sigma}_{v}\mu_{u}^{\top}\tilde{e}_{v}+\widetilde{\sigma}_{u}\mu_{v}^{\top}\tilde{e}_{u}\right)$ $=l_{0,uv}(\rho)R_{uv}+d_{uv}(\rho)R_{uv},$
for $R_{uv}\sim N(0,1)$ such that $R_{uv}$ is independent of $\mathbf{Y},\widetilde{\sigma}_{u}^{2},$ and $\widetilde{\sigma}_{v}^{2}$, with
$l_{0,uv}^{2}(\rho)=\rho^{2}\left(\sigma_{0v}^{2}\|\lambda_{0u}\|_{2}^{2}+\sigma_{0u}^{2}\|\lambda_{0v}\|_{2}^{2}\right)$
and
[…]
We can thus express
$\sqrt{n}(\widetilde{L}_{uv}-\mu_{u}^{\top}\mu_{v})=\dfrac{\sqrt{n}}{\sqrt{n+\tau^{-2}}}l_{0,uv}(\rho)R_{uv}+d_{uv}^{*}(\rho),$